Opened 7 years ago

Last modified 3 years ago

#364 assigned flight processing

EU10/03, flight day 234c/2010, Lake Balaton East Area A

Reported by: iopa Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/workspace/EU10_03-2010_234c_Lake_Balaton_East_Area_A

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Lake Balaton provides an ideal area for testing emerging algorithms on retrieval of in-water constituents and on mapping bloom-forming cyanobacteria using in-situ radiometric, airborne and satellite data.

Priority: a5

PI: Andrew Tyler

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

Sensors:

  • Eagle (27/10/2010)
  • Hawk (04/01/2011)
  • Leica LIDAR (Requested but low priority)
  • Digital Photography (27/10/2010)

Change History (28)

comment:1 Changed 7 years ago by iopa

  • Description modified (diff)

comment:2 Changed 7 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Started eagle processing.

comment:3 Changed 7 years ago by iopa

Applanix and IPAS have been done. The applanix data currently in the workspace covers both this project and also 234b.
For IPAS, there is processed IPAS in 234a in the workspace that covers 234a,234b,234c.
The proc directory from there could be entirely copied over to 234c or just symlinked, your choice.

comment:4 Changed 7 years ago by iopa

Taken camera processing.

comment:5 Changed 7 years ago by iopa

Camera ready for checking.

All images were processed, the test images were taken earlier in the day.

comment:6 Changed 7 years ago by emca

Delivery checked camera imagery:
~airborne/workspace/EU10_03-2010_234c_Lake_Balaton_East_Area_A/delivery/20101008/EU10_03

All data looks good. Ready to deliver.

comment:7 Changed 7 years ago by knpa

Found Eagle and Hawk SCTs:

eagle hawk
1 0.03 -0.01
2 0.05 0.00
3 0.00 0.00
4 0.04 0.00
5 0.00 -0.01
6 -0.01 -0.02

comment:8 Changed 7 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:9 Changed 7 years ago by knpa

Eagle awaiting delivery check.

comment:10 Changed 7 years ago by iopa

Eagle has been delivery checked;
the only issue found was that fastQC gives a header-filesize error when opening e234c061.bil
but everything seems to be in order between the header values and the filesize.

Moved .tex and .cfg files out of the delivery directory.

Eagle is ready to deliver.

comment:11 Changed 7 years ago by emca

  • Description modified (diff)

Eagle and Camera dispatched to PI 27/10/2010

comment:12 Changed 6 years ago by benj

  • Description modified (diff)

comment:13 Changed 6 years ago by knpa

Started preparing hawk for delivery.

comment:14 Changed 6 years ago by knpa

Line six is coming out with strange data values in level 1 and level 3. The line looks looks similar to the lines that were processed before (i.e. it has the wrong bandset). Reprocessing did not help and can not find any reason.

comment:15 Changed 6 years ago by knpa

benj and I have decided that though line 6 looks strange, there is nothing obviously wrong with it and it seems to have processed correctly (it is different from the previously processed - incorrect bandset - line 6).

Will deliver but with a warning included in Readme.

comment:16 Changed 6 years ago by benj

Just to elaborate on that comment, looking at the new and old versions of line 6, there is a consistent offset in values from one to the other and bad pixels are removed in the old version (they're not in the new version because we decided it wasn't getting all of them) - if it was somehow picking up the old calibration they wouldn't be. The offset is comparable to that for line 5, as are the magnitudes.

In other words, it looks funny when we look at it in gtviewer, but I suspect this is down to scaling and possibly sun angle, the data itself looks fine to me.

comment:17 Changed 6 years ago by knpa

Hawk ready for delivery check.

comment:18 Changed 6 years ago by knpa

Recreating level 1's due to calibration error.

comment:19 Changed 6 years ago by knpa

Level 1's put in delivery, ready to be checked.

comment:20 Changed 6 years ago by iopa

Hawk delivery has been checked;

lev1's show significant underflows which is to be expected.

The data & delivery seem in order, ready to go.

comment:21 Changed 6 years ago by knpa

  • Description modified (diff)

Hawk delivered on 04/01/11 to Andrew Tyler, Stirling.

comment:22 Changed 5 years ago by besm

Regarding the Lidar delivery..

  • PI name is wrong on lidar readme pdf
  • line numbers in log sheet and readme did not match
  • height differences were wrong (and not thorough enough)

comment:23 Changed 5 years ago by adbe

  • Component changed from Processing: general to Archiving

Changes made to LiDAR delivery.

comment:24 Changed 5 years ago by adbe

Reprocessing Hyperspectral delivery created.

comment:25 Changed 5 years ago by knpa

  • Milestone set to 2010 data processing completion

Beginning hyperspectral delivery check.

comment:26 Changed 5 years ago by knpa

Finished DC.

All looks good apart from:

  • e01 has an SCT offset

Have ran zip_mapped.sh

comment:27 Changed 5 years ago by adbe

Corrected sct offset

comment:28 Changed 3 years ago by knpa

Merged and deleted workspace version

Note: See TracTickets for help on using tickets.